OpenStudy (grimsnightmare):

A 25.0 mL sample of gas is enclosed in a gas-tight syringe at 22oC. If the syringe is immersed in an ice bath, what is the new gas volume? I have gathered that; V1=25.0mL T1=22C +273=295K V2=??? T2=273? I need to know if my information is correct.

OpenStudy (micahm):

what did you have in mind

OpenStudy (grimsnightmare):

Well I'm trying to use the Charles Law. However I am a little confused about what the second temperature will be. It says that the syringe is put in a ice bucket, then I have to find the new volume. I took a guess at 273K.

OpenStudy (micahm):

and how did it turned out

OpenStudy (grimsnightmare):

I ended up getting 23.1355. (Now I will round) However, I am given options for the answer and this doesn't match any of them. I don't know what I'm missing from the problem.

OpenStudy (micahm):

what grade is this

OpenStudy (grimsnightmare):

Well it's Chemistry 1

OpenStudy (micahm):

ok what grade is that

OpenStudy (grimsnightmare):

11

OpenStudy (grimsnightmare):

also 10th

OpenStudy (micahm):

ok i am not good in chem so much so here is some tips you can look up you question in google or @Preetha can maby help

OpenStudy (grimsnightmare):

Alright thanks

OpenStudy (micahm):

sorry i could not help

OpenStudy (grimsnightmare):

It's no problem. Thanks for attempting the problem and referring someone else.

OpenStudy (micahm):

@CaseyCarns

OpenStudy (grimsnightmare):

Thanks

OpenStudy (jfraser):

@grimsnightmare how does your equation look? What is the general form of Charles' Law?

OpenStudy (grimsnightmare):

sorry the general formula is V1/T1=V2/T2 However I changed it to: V2 = T2*V1/T1

OpenStudy (jfraser):

your rearrangement is correct, I don't know why you'd get an answer different than one of the options, unless there's a mistake in the problem itself. I get 23.15mL as well, so I don't know what to tell you

OpenStudy (grimsnightmare):

I figures out the problem. So if you were to work the problem by the original formula writing like this; 25.0 / 295 it equals 0.0847 You would round this right? My teacher didn't round to .085 so the problem was coming out all wrong.